Diffusion Morning Report 05/26/05

Ted Berg tberg at snf.stanford.edu
Thu May 26 06:45:06 PDT 2005


Good Morning All,
    Looking OK this morning. Epi burn box going back together. Tylan 1 
is still down for comm problem. Ebeam is down for cryo problems-
replacing relays for cold heads. Observing Nitride.

>   	Date: 	5/26/2005 	THICKNESS-Change at
> Tube Name/Number 	Status 	Comments 	 
> Tylan 1 	Down 	Comm failure 	 
> Tylan 2 	Up 	  	 
> Tylan 3 	Up 	  	 
> Tylan 4 	Up 	  	 
> Tylan 5-BBr3 	Up 	  	 
> Tylan 6-POCl3 	Up 	  	 
> Tylan 7-TEOS2 	Up 	  	 
> Tylan 8-SiGe 	Up 	  	3.12/30
> Tylan 9-Poly 	Up 	  	1.67/30
> Tylan 10-Nitride 	yellow 	Particles look good 	13.05/30
> Tylan 11-BPSG-LTO 	Up 	  	22.53/30
> Tylan 12-FGA 	Up 	  	 
> Tystar 	Up 	  	26.55/30
> Doped Poly 	GONE 	  	 
> TEOS 	yellow 	  	 
> Tylan 16 	N/A 	  	 
> Thermco 1 	Up 	  	 
> Thermco 2 	Up 	  	 
> Thermco Poly 	yellow 	testing doped 	 
> Thermco Nitride 	Down 	Waiting 	 
>   	  	  	 
> Burn Box 	Up 	  	 
> ZYGO 	up 	  	.
>   	  	  	 
> GAAS 22 	UP 	  	 
> GAAS 23 	UP 	  	 
>   	  	  	 
> SiH4 for furnaces 	  	B on line 345 PSI 	 changeover working well 
>

-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20050526/14cf6ab6/attachment.html>


More information about the diffusion mailing list